OpenStudy (anonymous):

Find the radius of the right cylinder shown, in which the height of the cylinder is equal to its diameter.

OpenStudy (anonymous):



OpenStudy (studygurl14):

What is S?

OpenStudy (anonymous):

Surface area.\[150\pi \]

OpenStudy (anonymous):

@studygurl14

geerky42 (geerky42):

\[ S=Ch+2\pi r^2, \text{ where }C\text{ is circumference of circle.} \\~\\150\pi=(2\pi r)h+2\pi r^2 \\~\\\text{We are given that }h=2r\text{, so}\\150\pi = (2\pi r)(2r)+2\pi r^2\\~\\ 150\pi=4\pi r^2+2\pi r^2\\~\\\text{Now combine like terms}\\150\pi=6\pi r^2\\~\\\text{Divide both sides by }6\pi\\\dfrac{150\pi}{6\pi}=r^2\\~\\ 25=r^2\\~\\\text{Take square root of both sides}\\ \boxed{r = 5} \]